OpenStudy (anonymous):

A bag contains 5 yellow marbles, 7 pink marbles, and 3 purple marbles. Gordon selects a marble without looking. What is the probability that he selects a yellow marble or a purple marble?

OpenStudy (kirbykirby):

There are 15 marbles in total. P(yellow) = 5/15 = 1/3 P(purple) = 3/15 = 1/5 P(yellow OR purple) = P(yellow) + P(purple) ... the addition rule applies! = 1/3 + 1/5 = 8/15